For the circuit shown, find Vab, It, I, & Iz using mode voltages. I 12 le r ab 4 skr 2 5khz 48V 7-yiokr
Answer #1

Solution 11 Na a 12k tatowy Tjske Žake ob Ske 1 Tisoke +1H Na-4860 + Vat va =0 2 5tis 12-170 es . . o s1+1+t (=2420° T 2 5+15It = va- 448LD - Va-4810*=-5.67+1-861° 2 It= 5:67-1660 A = 5.9094L-16.364A · tees Vab = I x 12 KR (V=2R) = (1.666 4+1:66621°)thank you please upvote

God bless you